OpenStudy (anonymous):

please solve: (7 x 10^−8 m2) − (4 x 10−8 m2)

OpenStudy (anonymous):

Important things to note here: - Your units are the same for both expressions, so they can be 'factored out'. - The orders of magnitude are also the same for both expressions, so they can also be factored out. \(7 \times 10^{-8} m^2 - 4 \times 10^{-8}m^2\) \(= (7 - 4) \times 10^{-8}m^2\)

OpenStudy (anonymous):

3 x 10^-8 (mass squared)